OpenStudy (anonymous):

The perimeter of a rectangle is 96 ft. Find the dimensions of the rectangle if the ratio of the length to the width is 7: 5. A. 30 ft by 18 ft B. 28 ft by 20 ft C. 24 ft by 24 ft D. 22 ft by 26 ft

OpenStudy (anonymous):

We have a Proportion ! we have : perimeter 2w+2L so we divided 96 on 2 have : 48 so W+L=48 And then : 7_____________x 5_______________y 12_______48 What happen ? 48 * 4 so we *4 to 5 and 7 : 4*7=28 4*5=20 Got it ?:) @Cookie_009